OpenStudy (anonymous):

Simplify: –4(x – 6) + 3(a – 7) –4x + 3a – 45 –4x + 3a + 45 –4x + 3a – 3 –4x + 3a + 3

OpenStudy (anonymous):

–4(x – 6) + 3(a – 7) = -4x + 24 + 3a - 21 = -4x + 3a + (24 - 21) = -4x + 3a + 3

OpenStudy (anonymous):

get it?